OpenStudy (anonymous):

(7-square root of 5)(7+5 square root of 5) find product .

OpenStudy (anonymous):

Here we're going to FOIL. \[(7-\sqrt{5})(7+5\sqrt{5})\] You can solve just as if it were binomials with variables.

OpenStudy (anonymous):

My options are 54 , 1225 , 7(square root of 5) and 44 .

OpenStudy (anonymous):

Let's go through the multiplication. First - would be 7*7 = 49 Outside - would be 7*5sqrt5 = 45sqt5 (think of the sqrt 5 as a variable like x) Inside - would be -sqrt5*7 = -7sqrt5 Last - would be -sqrt5*5sqrt5 = -25 Put that all together we get 49 + 45sqrt5 - 7sqrt 5 - 25 = ?